# Wasserstein Wormhole: Scalable Optimal Transport Distance with Transformers

### v1

@@ 6 Experiments
We demonstrate Wasserstein Wormhole on datasets representing multiple contexts where OT is frequently applied. We applied Wormhole to 2D point clouds derived from the MNIST and Fashion-MNIST datasets, and to 3D CAD models from ModelNet40 and ShapeNet [42]. As a high-dimensional, real-world application, we used Wormhole to embed multicellular microenvironments [43] from two single-cell-resolution spatial transcriptomics datasets [44, 45], along with groups of cells (’MetaCells’ [46]) from a scRNA-seq atlas of patient response to COVID [47]).
@@
We benchmarked our Wormhole against DiffusionEMD and Deep Wasserstein Embedding (DWE), both sharing Wormhole’s goal of finding OT preserving embeddings. We also compare against a constructed baseline approach we dub Fréchet [48] in which we use the closed-form $W_{2}$ distance for Gaussian distributions. For each point cloud, we calculated the mean ($\mu$) and covariance matrix ($\Sigma$) of its samples and use the following formula to approximate the OT distance between each pair of point clouds:
$W_{2}^{2}(X_{i},X_{j})\approx\|\mu_{i}-\mu_{j}\|^{2}+\text{Tr}\left(\Sigma_{i}+\Sigma_{j}-2\sqrt{\Sigma_{i}\Sigma_{j}}\right)$ (10)
The following sections present vignettes of Wormhole from several dataset, and results are summarized in Table 1. Wormhole outperforms all competing approaches and uniquely scales to massive & high-dimensional datasets such as the the 256,058-sample, 254-dimensional MERFISH Cell Niches cohort. We specifically highlight the improvement of Wormhole over DWE, which similarly learns a neural network parametric embedding. The transformer backbone of Wormhole proves to be superior to DWE’s CNN mode in low-dimensional dataset, while also essential in scaling Wormhole beyond 3D data.
@@ 6.1 MNIST (Touchstone)
We applied Wormhole to the MNIST dataset of hand-written digits, which consists of 70,000 samples of $28\times 28$ pixel grayscale images. We converted each sample to a point cloud by thresholding and selecting the set of pixel coordinates that passed the threshold (Figure 3a). While MNIST is considered small by modern machine-learning standards, straightforward computation of all $\binom{70000}{2}=2.45\cdot 10^{9}$ pairwise Wasserstein distances using current OT solvers is far from computationally feasible and would require a week of JAX-OTT [49] on an 80GB GPU (Figure 2).
@@ 6.3 ModelNet40 (Variational Wasserstein)
Fig. 5: Wormhole Variational Wasserstein. a. Mean encodings of each class were passed through the decoder to visualize their Wasserstein barycenters. Working in Wormhole space provides a quick and straightforward proxy for OT analysis, and recovered decodings reproduce the underlying class well. For brevity, we show 10 out of the 40 classes in from ModelNet40. b. Wormhole interpolation between two point clouds within the same class (top) and across different classes (bottom). Intermediate point clouds are linear interpolations between the embedding of each sample.
@@
Using Wormhole, we can also smoothly transition between point clouds via Wasserstein interpolation in the embedded space. Interpolating between distributions in Wasserstein space involves solving a weighted version of the barycenter problem as many times as there are intermediate samples. With Wormhole, we can leverage the Euclidean approximation of its latent space and directly decode linear interpolations between embeddings (Figure 5b). Despite the fact that interpolated embeddings are outside of the Wormhole training distribution, their decodings are valid and reliably visualize point cloud transitions.
6.4 ShapeNet (Gromov-
Wasserstein Wormhole)
Gromov-Wasserstein (GW) [50] is an extension of OT to distributions in different spaces that seeks the most isometric map between point clouds. Given two point clouds, their GW distance is:
@@ 6.4 ShapeNet (Gromov-
Where $d^{X}_{i,i^{\prime}}$ is the distance between $X_{i}$ and $X_{i^{\prime}}$, and similarly for $d^{Y}_{j,j^{\prime}}$. This problem is quadratic and thus, it is significantly more computationally intensive than standard OT [2]. GW has been a promising tool for aligning datasets from different modalities [51], mapping between graphs [52] and quantifying distances between 3D shapes with arbitrary orientation [53]. However, the utility of GW is curtailed by its complexity, motivating our Wormhole approach.
ShapeNet is a CAD-based dataset of 3D point clouds consisting of 16 different object classes. To construct the conditions where GW is usually applied, we randomly rotated each sample in 3D dimensions (Figure 6a). We then trained Wormhole to encode the rotated point clouds such that distances between embeddings retain GW distances.
@@ 6.5 Cellular Niches (High-Dimensional Data)
Spatially resolved transcriptomics (ST) are a set of technologies which capture both RNA information and the spatial context of cells to gain insight into tissue organization. OT has been frequently applied [54, 55] to analyze results of ST and unearth the relationship between a cell’s phenotype and its microenvironment. For example, Spatial Optimal Transport (SPOT) [56] models the niche of each cell as the point cloud from the gene expression of every cell in its environment, and defines differences between niches using Wasserstein.
While OT is an alluring metric to quantify cellular niches, SPOT cannot scale to ST datasets of typical size. As experimental technologies mature, atlas-level spatial assays continue to grow in ubiquity and size. One instance is the $254$-gene MERFISH dataset of the motor cortex [44], which consists of 256,058 cells across 64 individual slices.
We summarize the $50\mu m$ radius around each cell in the MERFISH dataset to construct niche point clouds in $254$ dimensions. OT linearization methods such as DWE and DiffusionEMD assume that data is voxelized or discretized. While low-dimensional datasets can be manipulated to fit within their frameworks, the curse of dimensionality prevents adequate voxelization of this Cell Niches dataset. As commercial technologies capture more genes [57], the need for a method that can scale to higher dimensions, such as Wormhole, becomes more acute.
Parvalbumin-positive (Pvalb) neurons are one of four major groups of inhibitory neuron, and are crucial for the regulation of neural circuitry. They are sparse in the motor cortex, comprising only $3\%$ of the dataset. To understand their spatial organization, it is therefore critical to study the entire cohort and integrate insights from multiple slices.
@@
After training Wormhole on the complete datasets (Table 1), we analyze the embeddings of the niches of Pvalb neurons. The first diffusion component (DC) [58] of Wormhole encodings aligns with their cortical depth across all slices within the dataset (Figure 7a-b). Despite the size and dimensionality of this dataset, Wormhole can still identify key biological signals across the cohort.
@@ 7 Discussion
We have introduced an approach, Wasserstein Wormhole, for embedding point clouds such that OT distances are preserved and can be approximated efficiently. Although OT is a popular approach, applying it to cohorts of distributions requires costly computation that scales quadratically with cohort size. Wormhole expands the functionality of OT to large datasets.
Encoding based on OT distances is theoretically challenged by the fact that Wasserstein is a non-Euclidean metric. To alleviate this concern, we derived theoretical lower and upper bounds that delimit how far non-Euclidean distance matrices can be from the embeddings. Our bounds grow smaller and tighter as the distance matrix approaches Euclidean. We further describe an algorithm based on projected gradient descent which is guaranteed to find an optimal embedding. On examples where the full Wasserstein distance matrix is tractable, Wormhole falls within the bounds and is close to the optimal error.
We applied our approach to six different settings, ranging from 2D point clouds to massive spatial transcriptomics datasets with hundreds of genes. Across all settings, Wormhole embeddings recapitulated the Wasserstein space, accurately encoded key aspects of the point clouds and outperformed competing approaches. Furthermore, Wormhole’s decoder recovered solutions to variational Wasserstein problems such as barycenters and point cloud interpolations. Wormhole seamlessly extends to other OT metrics, such as GW, and learns from arbitrary orientations.
Training Wasserstein Wormhole never exceeded 30 min on a single GPU, even on the 2048 points-per-sample ModelNet40 & ShapeNet datasets. Our algorithm is implemented in JAX [59] and integrates with OT tools (OTT-JAX) [49] for efficient Wasserstein distance calculations. Wormhole can be easily applied to any point cloud dataset via a pip installable package [^1] See tutorial in github.com/dpeerlab/WassersteinWormhole.
@@
We envision several future directions for this work. The complexity of transformers is quadratic in input size, limiting their scalability to point clouds with more than a few thousand samples. Alternatively, recently proposed models [60] can efficiently learn from point clouds by leveraging local geometry. Fitting those models within the Wormhole framework could produce a more versatile approach.

### v4

@@ 6 Experiments
We demonstrate Wasserstein Wormhole on datasets representing multiple contexts where OT is frequently applied. We applied Wormhole to 2D point clouds derived from the MNIST and Fashion-MNIST datasets, and to 3D CAD models from ModelNet40 and ShapeNet (Chang et al., 2015). As a high-dimensional applications, we used Wormhole to embed multicellular niches (Haviv et al., 2023) from two high-resolution spatial transcriptomics datasets (Zhang et al., 2021; Lohoff et al., 2022), along with groups of cells (’MetaCells’ (Persad et al., 2023)) from a scRNA-seq atlas of patient response to COVID (Stephenson et al., 2021).
We benchmarked our Wormhole against DiffusionEMD and Deep Wasserstein Embedding (DWE), both sharing Wormhole’s goal of finding OT preserving embeddings. We also compare against a constructed baseline we dub Fréchet (Dowson & Landau, 1982) in which we use the closed-form $W_{2}$ distance for Gaussian distributions. For each point cloud, we calculated the mean ($\mu$) and covariance matrix ($\Sigma$) of its samples and use the following formula to approximate the OT distance between each point cloud pair:
$W_{2}^{2}\approx\|\mu_{i}-\mu_{j}\|^{2}+\text{Tr}\left(\Sigma_{i}+\Sigma_{j}-2\sqrt{\Sigma_{i}\Sigma_{j}}\right)$ (10)
The following sections present vignettes of Wormhole from several dataset, and results are summarized in Table 1. Wormhole outperforms all competing approaches and uniquely scales to massive & high-dimensional datasets such as the the 256,058-sample, 254-dimensional MERFISH Cell Niches cohort. We specifically highlight the improvement of Wormhole over DWE, which similarly learns a neural network parametric embedding. The transformer backbone of Wormhole is superior to DWE’s CNN in low-dimensional datasets, while also proving to be essential in scaling Wormhole beyond 3D data.
@@ 6.1 MNIST (Touchstone)
We applied Wormhole to the MNIST dataset of hand-written digits, which consists of 70,000 samples of $28\times 28$ pixel grayscale images. We converted each sample to a point cloud by thresholding and selecting the set of pixel coordinates that passed the threshold (Figure 3a). While MNIST is considered small by modern machine-learning standards, straightforward computation of all $\binom{70000}{2}=2.45\cdot 10^{9}$ pairwise Wasserstein distances using current OT solvers is far from computationally feasible and would require a week of JAX-OTT (Cuturi et al., 2022) on an 80GB GPU (Figure 2).
@@ 6.3 ModelNet40 (Variational Wasserstein)
Fig. 5: Wormhole Variational Wasserstein. a. Mean encodings of each class were passed through the decoder to visualize their Wasserstein barycenters. Working in Wormhole space provides a quick and straightforward proxy for OT analysis, and recovered decodings reproduce the underlying class well. For brevity, we show 10 out of the 40 classes in from ModelNet40. b. Wormhole interpolation between two point clouds within the same class (top) and across different classes (bottom). Intermediate point clouds are decodings of linear interpolations between embeddings.
@@
Using Wormhole, we can also smoothly transition between point clouds via Wasserstein interpolation in the embedded space. Interpolating between distributions in Wasserstein space involves solving a weighted version of the barycenter problem as many times as there are intermediate samples. With Wormhole, we can leverage the Euclidean approximation of its latent space and directly decode linear interpolations of embeddings (Figure 5b). Despite interpolated embeddings lying outside of Wormhole training distribution, their decodings reliably visualize point cloud transitions.
6.4 ShapeNet (Gromov-Wasserstein Wormhole)
Gromov-Wasserstein (GW) (Mémoli, 2011) is an extension of OT which seeks the most isometric map between point clouds. Given two point clouds, their GW distance is:
@@ 6.4 ShapeNet (Gromov-Wasserstein Wormhole)
Where $d^{X}_{i,i^{\prime}}$ is the distance between $X_{i}$ and $X_{i^{\prime}}$, and similarly for $d^{Y}_{j,j^{\prime}}$. This problem is quadratic is significantly more computationally intensive than standard OT (Peyré et al., 2019). Unlike standard OT, GW can align datasets from different modalities (Demetci et al., 2020), map between graphs (Xu et al., 2019) and quantify distances between 3D shapes with arbitrary orientation (Mémoli, 2009).
ShapeNet is a CAD-based dataset of 3D point clouds consisting of 16 different object classes. To construct the conditions where GW is usually applied, we randomly rotated each sample in 3D dimensions (Figure 6a). Standard OT is blind to structure when orientation is unknown, making rotation-invariant GW necessary. We then trained Wormhole to encode the rotated point clouds such that distances between embeddings retain GW distances.
@@ 6.5 Cellular Niches (High-Dimensional Data)
Spatially resolved transcriptomics (ST) are a set of technologies which capture both RNA information and the spatial context of cells to gain insight into tissue organization. OT has been frequently applied to analyze results of ST and unearth the relationship between a cell’s phenotype and its microenvironment (Cang & Nie, 2020; Haviv et al., 2024). For example, Spatial Optimal Transport (SPOT) (Mani et al., 2022) models the niche of each cell as the point cloud from the gene expression of every cell in its environment, and defines differences between niches using Wasserstein.
While OT is an alluring metric to quantify cellular niches, SPOT cannot scale to ST datasets of typical size. As experimental technologies mature, atlas-level spatial assays continue to grow in ubiquity and size. For example, the $254$-gene MERFISH dataset of the motor cortex (Zhang et al., 2021), which consists of 256,058 cells across 64 slices.
We summarize the $50\mu m$ radius around each cell in the MERFISH dataset to construct niche point clouds in $254$ dimensions. OT linearization methods such as DWE and DiffusionEMD assume that data is voxelized or discretized. While low-dimensional datasets can be manipulated to fit within their frameworks, the curse of dimensionality prevents adequate voxelization of this Cell Niches dataset. As commercial technologies capture more genes (Marco Salas et al., 2023), the need for a method that can scale to higher dimensions, such as Wormhole, becomes more acute.
@@
Parvalbumin-positive (Pvalb) interneurons are one the major groups of inhibitory neuron in the cortex, and are crucial for the regulation of neural circuitry (Sohal et al., 2009). They are sparse in the motor cortex, comprising only $3\%$ of the dataset. To understand their spatial organization, it is therefore critical to study the entire cohort and integrate insights from multiple tissue slices.
After training Wormhole on the complete datasets (Table 1), we analyze the embeddings of the niches of Pvalb neurons. The first diffusion component (DC) (Coifman & Lafon, 2006) of Wormhole encodings aligns with their cortical depth across all slices within the dataset (Figure 7a-b). Despite the size and dimensionality of this dataset, Wormhole can still identify key biological signals across the cohort.
@@ 7 Discussion
We have introduced Wasserstein Wormhole, an approach for embedding point clouds such that OT distances are preserved and can be approximated efficiently. Although OT is ubiquitous, applying it to cohorts of distributions requires costly computations and scales quadratically with cohort size. Wormhole expands OT to large datasets.
Encoding based on OT distances is theoretically challenged by the fact that Wasserstein is a non-Euclidean metric. To alleviate this concern, we derived theoretical lower and upper bounds that delimit how far non-Euclidean distance matrices can be from the embeddings. We further describe an algorithm based on projected gradient descent which is guaranteed to find an optimal embedding. On examples where the full Wasserstein distance matrix is tractable, Wormhole falls within the bounds and is close to the optimal error.
We applied our approach to eight different settings, ranging from 2D point clouds to massive spatial transcriptomics datasets with hundreds of genes. Across all settings, Wormhole embeddings recapitulated the Wasserstein space, accurately encoded key aspects of the point clouds and outperformed competing approaches. Furthermore, Wormhole’s decoder recovered solutions to variational Wasserstein problems such as barycenters and point cloud interpolations. Wormhole seamlessly extends to other OT metrics, such as GW, and learns from arbitrary orientations.
Training Wasserstein Wormhole never exceeded 30 min on a single GPU, even on the 2048 points-per-sample ModelNet40 & ShapeNet datasets. Our algorithm is implemented in JAX (Bradbury et al., 2021) and integrates with OT tools (OTT-JAX) (Cuturi et al., 2022) for efficient Wasserstein distance calculations. Wormhole can be easily applied to any point cloud dataset via a pip installable package.
@@
We envision several future directions for this work. Currently, Wormhole is only equipped to handle uniform point-clouds, where each point is equally weighted. Future work can extend the Attention mechanism to include arbitrarily weighted point-clouds, to reflect more general instances of empirical distributions and applications of OT.
@@
Acknowledgements
This work was supported by National Cancer Institute grants P30 CA08748 and U54 CA209975. Dana Pe’er is an Investigator at the Howard Hughes Medical Institute.
Impact Statement
Wasserstein Wormhole is a novel method to scale computations of OT distances to large cohorts. We currently do not foresee any adverse effects of our work